# Towards Faithful and Plausible Natural Language Explanations for Image Classification: A Pipeline Approach

### v1

@@ Appendix G Details on stability experiments
    inter-set stability: We compare pairs of unaltered images ($1^{st}$ set of images) and corresponding images mapped by the $\zeta$ function ($2^{nd}$ set of images). After the $\zeta$ mapping is performed, we run the proposed method on both images, yielding two explanations. We then compute various language similarity metrics between the two explanations.
@@
    intra-set stability: We compare explanations pairs produced for different images to verify the diversity of generated explanations.
The computations were conducted on explanations produced by the proposed method for a 500-element subset of validation ImageNet data. The subset was constructed by randomly picking 50 examples from 10 selected, diverse classes (library, over skirt, palace, prison, wall clock, lakeside, coral reef, volcano, fountain, basset).
@@
Appendix H Details on the covering experiment
@@
Based on the following explanation of the classifier’s prediction, cover the reasons for its decision with white rectangles. You can use as many rectangles as you like, but the total area of the covered image cannot be larger than half of the image. If it is not possible to cover the mentioned reasons by converting only 50% of the image, please do your best to cover the most important information.
@@
Appendix I Layer-wise Relevance Propagation
@@
First, unlike many other methods, LRP works at the neuron level, which is strictly required by our method. Therefore, methods that provide pixel-level importance scores such as RISE Petsiuk et al. (2018) or Grad-CAM Selvaraju et al. (2017b), methods that typically work on image segments such as LIME Ribeiro et al. (2016b) or SHAP Lundberg and Lee (2017) are not suitable for our approach.
@@
Appendix J Statistical analysis of the human evaluation results

### v2

@@ Appendix F Details on stability experiments
    intra-set stability: We compare pairs of unaltered images ($1^{st}$ set of images) and corresponding images mapped by the $\zeta$ function ($2^{nd}$ set of images). After the $\zeta$ mapping is performed, we run the proposed method on both images, yielding two explanations. We then compute various language similarity metrics between the two explanations.
@@
    inter-set stability: We compare explanations pairs produced for different images to verify the diversity of generated explanations.
The computations were conducted on explanations produced by the proposed method for a 500-element subset of validation ImageNet data. The subset was constructed by randomly picking 50 examples from 10 selected, diverse classes (library, over skirt, palace, prison, wall clock, lakeside, coral reef, volcano, fountain, basset) in a stratified manner.
@@
Appendix G Details on the covering experiment
@@
Based on the following explanation of the classifier’s prediction, cover the reasons for its decision with white rectangles. You can use as many rectangles as you like, but the total area of the covered image cannot be larger than half of the image. If it is not possible to cover the mentioned reasons by covering only 50% of the image, please do your best to cover the most important information.
@@
Appendix H Layer-wise Relevance Propagation
@@
First, unlike many other methods, LRP works at the neuron level, which is strictly required by our method. Therefore, methods that provide pixel-level importance scores such as RISE Petsiuk et al. (2018) or Grad-CAM Selvaraju et al. (2017), methods that typically work on image segments such as LIME Ribeiro et al. (2016) or SHAP Lundberg and Lee (2017) are not suitable for our approach.
@@
Appendix I Statistical analysis of the human evaluation results